OpenStudy (anonymous):

help!!! will fan & medal (: An oil company fills 1 over 12 of a tank in 1 over 3 hour. At this rate, which expression can be used to determine how long will it take for the tank to fill completely? 1 over 12 ⋅ 3 hours 1 over 3 ⋅ 12 hours 1 over 3 ⋅ 1 over 12 hour 3 ⋅ 12 hours

OpenStudy (anonymous):

x = time to fill a whole tank \[\frac{ \frac{ 1 }{ 12 } }{ \frac{ 1 }{ 3 } }=\frac{ 1 }{ x }\]

OpenStudy (anonymous):

1/3 * 12 hours because if you think about it, 1/12 of a tank only takes 1/3 of time to fill, so you can then multiply 1/3 by 12 to get how long it will take to fill 1 tank.

OpenStudy (anonymous):

which is 4 hours

OpenStudy (anonymous):

ohhhhhhhhhh ok, thank you both!!!!